OpenStudy (anonymous):

Solve for x if: 2(x+2)=10

OpenStudy (anonymous):

x is 3

OpenStudy (fuzzbuzz):

x+2 = 10/2 therefore, x=?

OpenStudy (anonymous):

thanks